# Emotion-Coherent Reasoning for Multimodal LLMs via Emotional Rationale Verifier

### v2

@@ 1 Introduction
As multimodal large language models (MLLMs) (Chen et al. 2024; Li et al. 2024; Wang et al. 2024; Chu et al. 2024; Cheng et al. 2024b) continue to evolve, they are reshaping human-computer interaction (HCI) by enabling systems to interpret cross-modal inputs (face video and speech audio), perform contextual reasoning, and generate human-like responses (Fei et al. 2024; Park et al. 2024; Kim et al. 2025). To support truly natural and emotionally aware communication, such systems must go beyond surface-level responses and capture the nuance, intent, and emotional context of user inputs. Among these capabilities, emotion understanding plays a central role in facilitating deeper, more meaningful interaction—particularly in applications such as psychological counseling (Schmidgall et al. 2024; Lee et al. 2024), educational guidance (Schutz and Pekrun 2007), and empathetic dialogue systems (Fei et al. 2024; Zhang et al. 2025).
@@
While conventional multimodal emotion recognition (MER) approaches have treated emotion understanding as a classification task (Zhang et al. 2024b), single-label predictions often fail to capture the complexity of human affect conveyed through face video and speech audio. To address this limitation, recent work has turned to descriptive emotion understanding, leveraging MLLMs to generate natural language explanations that justify predicted emotions (Lian et al. 2023; Cheng et al. 2024a; Lian et al. ; Yang et al. 2025a). This shift is especially valuable in HCI scenarios, where users not only expect accurate emotional classification from visual and auditory cues but also seek coherent and human-like justifications to build trust and ensure transparency.
@@
To enhance emotion reasoning, a recent study leverages Reinforcement Learning (RL) and Chain-of-Thought (CoT) prompting to train MLLMs (Zhao, Wei, and Bo 2025). Although this method improves classification accuracy, it fails to consistently produce emotionally coherent explanations. This limitation may be attributable to the nature of the datasets used in current training pipelines. While recent method leverages a variety of emotion datasets, such as EMER (Lian et al. 2023), MERR (Cheng et al. 2024a), and MER-Caption (Lian et al. 2025), only a small fraction of these include human-annotated rationales. The majority of training relies on larger datasets without explanation labels, where reward signals focus solely on the correctness of predicted emotions, regardless of whether the generated rationale is emotionally coherent. As a result, models may generate emotionally incoherent explanations without penalty, leading to a misalignment between predictions and rationales, as shown in Figure 1.
@@
In this paper, we propose a novel framework for generating emotionally coherent explanations, even when trained on emotion datasets lacking human-annotated rationales. Our approach is based on the key assumption that LLMs possess sufficient linguistic and emotional knowledge to evaluate the consistency between a given emotion and a generated explanation. However, incorporating LLMs directly into the evaluation process can be computationally expensive and impractical in deployment settings. To address this, we distill the knowledge of the LLM into a lightweight model, referred to as the Emotional Rationale Verifier (ERV), which efficiently assesses the alignment between an explanation and the target emotion. Building on this verifier, we integrate ERV into a reward-driven training pipeline. In this setup, the model receives an explanation reward based on how well its generated explanation aligns with the GT emotion as judged by ERV. To further improve reward fidelity, we decompose each explanation into individual sentences, filter out emotionally neutral ones (e.g., those describing background or appearance), and assign rewards based on the proportion of emotionally salient sentences that correctly reflect the target emotion.
@@
Furthermore, to evaluate the emotional coherence, we introduce a set of novel metrics: Explanation Emotion Accuracy (EEA), which measures alignment between the explanation and the target emotion, Explanation–Prediction Consistency (EPC), which quantifies the internal consistency between the explanation and predicted emotion, and Faithful Consistency Rate (FCR), a strict metric that requires all three components, explanation, predicted emotion, and target emotion, to agree. These metrics provide deeper insight into the emotional coherence of generated explanations beyond standard classification accuracy.
@@ 2 Related Works
2.1 Mutlmodal Emotion Recognition (MER)
The task of Multimodal Emotion Recognition (MER) aims to understand human emotional states by analyzing diverse multimodal cues present in expressive scenes. From audio and visual cues to text, body, and physiological signals, numerous studies (Kossaifi et al. 2017; Zadeh et al. 2018; Poria et al. 2018; Busso et al. 2008; Perepelkina, Kazimirova, and Konstantinova 2018; Liu et al. 2022; Jiang et al. 2020; Livingstone and Russo 2018; Lee et al. 2019; Wang et al. 2022) have proposed datasets consisting of emotion-related modalities and corresponding emotion labels. With those datasets, traditional approaches have primarily relied on modality-specific encoders (Zhang et al. 2017, 2021; Zhao and Liu 2021; Sun et al. 2023; Chumachenko, Iosifidis, and Gabbouj 2024) for feature extraction and fusion strategies (Narayan et al. 2024; Sun et al. 2024), achieving solid performance in emotion classification but falling short in terms of explainability.
However, unlike clear object classification, emotion recognition suffers from inherent ambiguity, as emotion labels alone often fail to fully capture an individual’s state. To address the challenge of unreliable labeling caused by the limited expressivity of fixed label mappings, multimodal emotion descriptive dataset EMER (Lian et al. 2023) has been proposed. This dataset not only requires models to predict emotion labels from visual, audio, and textual cues in videos, but also to generate descriptive explanations of the emotional states. Additionally, with the rise of Large Language Models (LLMs) (Touvron et al. 2023; Team 2024; Grattafiori et al. 2024), leveraging the natural language generation power, recent instruction-tuned multimodal LLMs approaches (Xing et al. 2024; Cheng et al. 2024a; Lian et al. ; Yang et al. 2025a; Hu et al. 2025; Zhao et al. 2025a) move beyond classification toward emotion reasoning to enable interpretable, context-aware emotion understanding.
However, as these descriptions are based on human annotations, dataset sizes are limited, and scaling up emotional descriptive datasets is constrained by annotation costs. To address this issue, recent studies (Cheng et al. 2024a; Lian et al. 2024; Yang et al. 2025a) have investigated using the generative power of MLLMs to automatically augment emotional descriptive datasets. Building on these description-rich emotion datasets, many studies (Cheng et al. 2024a; Lian et al. 2024) have explored interpreting and generating explanations grounded in video and audio cues from samples. Furthermore, recent architectural advancements (Zhao et al. 2025b; Lian et al. ) have contributed to improved recognition accuracy. However, these approaches typically rely on supervised fine-tuning (SFT), which inherently limits generalization due to its dependence on the scale and diversity of annotated emotional descriptions.
@@ 2.2 GRPO-Based RL with Verifiable Rewards
Recent work has shown that RL post-training can substantially sharpen the CoT abilities of MLLMs. In particular, Group Relative Policy Optimization (GRPO) (Shao et al. 2024) combined with Verifiable Reward (Guo et al. 2025) achieves strong generalization even with limited data and fixed model backbones, excelling in math and coding benchmarks.
Motivated by this success, GRPO has been adapted to a wide range of multimodal tasks, including emotion recognition (Zhao, Wei, and Bo 2025), general vision tasks (Chen et al. 2025a), object detection (Liu et al. 2025; Shen et al. 2025; Park et al. 2025), and the video domain (Li et al. 2025; Lee et al. 2025; Xing et al. 2025), often designing reward functions tailored to their specific task objectives. However, these approaches predominantly focus on optimizing the final answer, paying little attention to the quality and consistency of the intermediate reasoning steps, a degradation problem highlighted by  (Wei et al. 2025). To address this, recent studies (Chen et al. 2025b; Yang et al. 2025b) have introduced consistency-aware methodologies to improve reasoning and answer alignment. Yet, those consistency problems have not yet been explored in inherently ambiguous domains such as MER. We therefore introduce a novel reward applied with GRPO training that explicitly enforces coherence between the generated explanation and the target emotion.
@@ 3 Method
Our primary objective is to develop a Multimodal Large Language Model (MLLM) capable of emotion understanding that not only accurately predicts emotion labels from face video and speech audio but also generates consistent and faithful natural language explanations aligned with these predictions. Building upon recent advancements demonstrating the effectiveness of Reinforcement Learning (RL)-based training for improving both emotion recognition accuracy and reasoning in MLLM (Zhao, Wei, and Bo 2025), we integrate a novel Emotional Rationale Verifier (ERV) module into an established RL framework. This integration allows us to directly evaluate and enhance the emotional consistency of the generated explanations. Specifically, we extend the HumanOmni model architecture (Zhao et al. 2025b) and Group Relative Policy Optimization (GRPO) (Guo et al. 2025), a reward-driven RL training strategy showing strong performance in various multimodal tasks (Liu et al. 2025; Chen et al. 2025a; Zhao, Wei, and Bo 2025). While HumanOmni effectively unifies multimodal features and GRPO enhances Chain-of-Thought (CoT) reasoning, existing approaches often struggle with emotionally incongruent explanations despite accurate label predictions, as highlighted in our introduction.
To precisely address this critical misalignment, our proposed ERV module assesses whether a generated explanation coherently supports the target emotion. Based on this assessment, we define an additional reward signal, the Explanation Reward ($R_{E}$), which encourages the MLLM to produce more faithful and emotion-aligned rationales during RL fine-tuning. The overall training pipeline, detailing the ERV module and its integration into the GRPO framework, is visually represented in Figure 2.
@@ 3.1 Model Architecture and Training Overview
The training of this model follows a robust two-stage strategy: an initial Supervised Fine-Tuning (SFT) (Zhang et al. 2024a) phase, followed by RL using GRPO. This progressive approach ensures the model first learns fundamental reasoning structures and then refines its explanation generation based on explicit reward signals.
Initializing Basic Emotion Reasoning via SFT
The initial training phase involves SFT on a high-quality, albeit relatively small, dataset named EMER (Lian et al. 2023). This dataset consists of $(V,A)$ pairs annotated with precise emotion labels and corresponding descriptive explanations. The primary objective of this phase is to enable the MLLM to grasp the basic structural relationships between multimodal inputs, emotion labels, and their textual explanations. It specifically helps the model learn to produce outputs in the required format <<think>><</think>><<answer>><</answer>>. This SFT phase serves as a crucial initialization step for our policy model, denoted as $\pi_{\theta}$ (i.e., the trainable MLLM). During SFT, the model is optimized to maximize the likelihood of generating the target output sequence given the input. The training objective is defined by the negative log-likelihood loss:
@@
Figure 2: (a) Enhancing explanation coherence using Explanation-Guided Reinforcement Learning (RL). The policy model $\pi_{\theta}$ generates $G$ responses $o_{1}$, …, $o_{G}$, and the Emotional Rationale Verifier (ERV) assigns an explanation reward $R_{E}$ to each response. (b) Explanation Reward $R_{E}$ Computation. For each response $o_{i}$, the corresponding explanation $E_{i}$ is extracted and evaluated to produce its reward $R_{i,E}$. Collectively, $R_{E}=\{R_{1,E},R_{2,E},\ldots,R_{G,E}\}$ denotes the set of explanation rewards for the $G$ generated responses.
@@
Following the SFT phase, we transition to RL using GRPO to enhance the model’s reasoning capability and, critically, the consistency of its explanations. The SFT-trained model $\pi_{\theta}$ functions as the learnable policy, while a frozen duplicate, $\pi_{ref}$, acts as the reference model. For each input $(V,A,P)$, both the policy model and the reference model generate G candidate output sequences, $\{o_{1},o_{2},\dots,o_{G}\}$. Each generated sequence is then evaluated by a comprehensive reward function, which we detail in the next section. The policy $\pi_{\theta}$ is subsequently optimized to maximize the expected reward $R$, while simultaneously being regularized to remain close to the behavior of $\pi_{ref}$. The optimization objective of GRPO is defined as follows:
@@
While the GRPO framework is highly effective for reward-driven optimization, it traditionally lacks an explicit mechanism to guarantee the consistency between the predicted emotion and its generated explanation. To overcome this inherent limitation and directly address the core challenge identified in our introduction, we introduce the Emotional Rationale Verifier ($ERV$) and its associated reward ($R_{E}$) as a novel, crucial signal within the GRPO training loop.
@@ 3.2 Emotional Rationale Verifier (ERV): Design and Training
Although existing datasets such as EMER (Lian et al. 2023) and MERR-fine (Cheng et al. 2024a) provide pairs of emotional descriptions and emotion labels, their limited sizes (332 and 4,487 samples, respectively) are insufficient for our setting. This is particularly problematic because the generated descriptions often contain not only emotional expressions but also contextual details about people, environments, and temporally evolving emotional states. In addition, the emotion label distribution is imbalanced, as shown in Appendix B.1. As a result, existing datasets lack the diversity and coverage necessary to effectively train the Emotional Rationale Verifier (ERV).
To address the limitations of existing datasets in terms of both size and class imbalance, we construct a pseudo-labeled dataset consisting of emotional text descriptions paired with emotion labels. As shown in Figure 1, while R1-Omni (Zhao, Wei, and Bo 2025) does not always produce perfectly emotion-aligned outputs, its generated descriptions are often emotionally plausible. Leveraging this property, we generate textual explanations from the training sets of DFEW (Jiang et al. 2020) and MAFW (Liu et al. 2022), and use a closed-source LLM (GPT-4.1) to assign up to two representative emotion labels to each description. This process yields a pseudo-labeled dataset of 20K emotional text–label pairs, which is used to train our ERV module. To further mitigate the class imbalance, we augment underrepresented categories, namely disgust, fear, contempt, disappointment, neutral, and helplessness, by generating additional emotional descriptions using GPT-4.1 (see Figure 7). We provide few-shot exemplars for each of these emotion labels as prompts, and synthesize new descriptions such that each of these categories contains approximately 1K examples in the final dataset.
To train the ERV on the constructed dataset, we use the RoBERTa model trained on the GoEmotions dataset (Demszky et al. 2020) [^1] https://huggingface.co/SamLowe/roberta-base-go˙emotions as a backbone. We fine-tune this model to predict emotion labels from text descriptions using our 23K-sample dataset, which includes both human-annotated and GPT-generated pseudo-labeled examples. The model predicts the emotion label based on the output of the [CLS] token, which encodes the global semantics of the input description. These predictions are then used to compute the Explanation Reward for each generated explanation $E_{i}$.

### v3

@@ Abstract
Codehttps://github.com/Rhatanii/ERV
Extended versionhttps://arxiv.org/abs/2510.23506
@@ 1 Introduction
As multimodal large language models (MLLMs) (chen2024internvl; li2024llava; wang2024qwen2vl; chu2024qwen2; cheng2024videollama) continue to evolve, they are reshaping human-computer interaction (HCI) by enabling systems to interpret cross-modal inputs (face video and speech audio), perform contextual reasoning, and generate human-like responses (fei2024empathyear; park2024let; kim2025speaking). To support truly natural and emotionally aware communication, such systems must go beyond surface-level responses and capture the nuance, intent, and emotional context of user inputs. Among these capabilities, emotion understanding plays a central role in facilitating deeper, more meaningful interaction—particularly in applications such as psychological counseling (schmidgall2024agentclinic; lee2024cactus), educational guidance (schutz2007emotion), and empathetic dialogue systems (fei2024empathyear; zhang2025towards).
@@
While conventional multimodal emotion recognition (MER) approaches have treated emotion understanding as a classification task (zhang2024traditionalmer), single-label predictions often fail to capture the complexity of human affect conveyed through face video and speech audio. To address this limitation, recent work has turned to descriptive emotion understanding, leveraging MLLMs to generate natural language explanations that justify predicted emotions (lian2023EMER; cheng2024MERR; lianaffectgpt25; yang2025omni). This shift is especially valuable in HCI scenarios, where users not only expect accurate emotional classification from visual and auditory cues but also seek coherent and human-like justifications to build trust and ensure transparency.
@@
To enhance emotion reasoning, a recent study leverages Reinforcement Learning (RL) and Chain-of-Thought (CoT) prompting to train MLLMs (zhao2025r1). Although this method improves classification accuracy, it fails to consistently produce emotionally coherent explanations. This limitation may be attributable to the datasets used in current training pipelines. While recent method leverages a variety of emotion datasets, such as EMER (lian2023EMER), MERR (cheng2024MERR), and MER-Caption (lianaffectgpt25), only a small fraction of these include human-annotated rationales. The majority of training relies on larger datasets without explanation labels, where reward signals focus solely on the correctness of predicted emotions, regardless of whether the generated rationale is emotionally coherent. As a result, models generate emotionally incoherent explanations without penalty, leading to a misalignment between predictions and rationales, as shown in Figure 1.
@@
In this paper, we propose a novel framework for generating emotionally coherent explanations, even when trained on emotion datasets lacking human-annotated rationales. Our approach is based on the key assumption that LLMs possess sufficient linguistic and emotional knowledge to evaluate the consistency between a given emotion and a generated explanation. However, incorporating LLMs directly into the evaluation process can be computationally expensive and impractical in deployment settings. To address this, we distill the knowledge of the LLM into a lightweight model, referred to as the Emotional Rationale Verifier (ERV), which efficiently assesses the alignment between an explanation and the target emotion. Building on this verifier, we integrate ERV into a reward-driven training pipeline. In this setup, the model receives an explanation reward based on how well its generated explanation aligns with the target emotion as judged by ERV. To further improve reward fidelity, we decompose each explanation into individual sentences, filter out emotionally neutral ones (e.g., those describing background or appearance), and assign rewards based on the proportion of emotionally salient sentences.
@@
Furthermore, to evaluate the emotional coherence, we introduce a set of novel metrics: Explanation Emotion Accuracy (EEA), which measures alignment between the explanation and the target emotion, Faithful Consistency Rate (FCR), requires all three components, explanation, predicted emotion, and target emotion, to agree, and Explanation–Prediction Consistency (EPC), which quantifies the consistency between the explanation and predicted emotion. These metrics provide deeper insight into the emotional coherence of generated explanations beyond standard classification accuracy.
@@ 2 Related Works
2.1 Multimodal Emotion Recognition
The task of Multimodal Emotion Recognition (MER) aims to understand human emotional states by analyzing diverse multimodal cues in expressive scenes. From audio and visual cues to text, body, and physiological signals, numerous studies (kossaifi2017afew; zadeh2018multimodal; poria2019meld; busso2008iemocap; perepelkina2018ramas; liu2022mafw; jiang2020dfew; livingstone2018RAVDESS; lee2019CAER; wang2022ferv39k) have proposed datasets consisting of emotion-related modalities and corresponding emotion labels. With those datasets, traditional approaches have primarily relied on modality-specific encoders (zhang2017speech; zhang2021learning; zhao2021former; sun2023mae; chumachenko2024mma) for feature extraction and fusion strategies (narayan2025facexformer; sun2024hicmae), achieving solid performance in emotion classification.
However, unlike object classification, emotion classification suffers from inherent ambiguity, as emotion labels alone often fail to fully capture an individual’s state. To address the challenge of unreliable labeling caused by the limited expressivity of fixed label mappings, multimodal emotion descriptive dataset EMER (lian2023EMER) has been proposed. This dataset requires models not only to predict emotion labels from visual, audio, and textual cues, but also to generate descriptive explanations of the emotional states. Additionally, with the rise of Large Language Models (LLMs) (touvron2023llama2; team2024qwen2; grattafiori2024llama), and their powerful natural language generation capabilities, recent instruction-tuned MLLM approaches (xing2024emo; cheng2024MERR; lianaffectgpt25; yang2025omni; hu2025feallm; zhao2025favchat) move beyond classification toward reasoning, enabling interpretable and context-aware emotion understanding.
Furthermore, motivated by the increasing use of rich emotional descriptions, recent studies (cheng2024MERR; lian2024open; yang2025omni) have explored leveraging the generative capabilities of LLMs to automatically expand emotion descriptive datasets. These approaches aim to enrich the diversity and expressivity of descriptive samples by generating additional explanations grounded in multimodal cues. Building on such augmented and description-rich emotion datasets, many works (cheng2024MERR; lian2024open) have investigated interpretable reasoning over video and audio, while architectural advancements (zhao2025humanomni; lianaffectgpt25) have further improved recognition accuracy. However, most existing methods rely on supervised fine-tuning (SFT), which limits generalization due to its dependence on the scale and diversity of annotated emotional descriptions.
@@ 2.2 GRPO-Based RL with Verifiable Rewards
Recent work has shown that RL post-training can substantially sharpen the CoT abilities of MLLMs. In particular, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) combined with Verifiable Reward (guo2025deepseek) achieves strong generalization with limited data and fixed model backbones, excelling in math and coding benchmarks.
Motivated by this success, GRPO has been adapted to a wide range of multimodal tasks, including emotion recognition (zhao2025r1), general vision tasks (chen2025r1v), object detection (liu2025visualrft; shen2025vlm; park2025dip), and the video domain (li2025videochat; lee2025refocus; xing2025echoink), often designing reward functions tailored to their specific task objectives. However, these approaches predominantly focus on optimizing the final answer, paying little attention to the quality and consistency of the intermediate reasoning steps, a degradation problem highlighted by  (wei2025gtr).
To address this, recent studies (chen2025grpocare; yang2025humanomniv2) have introduced consistency-aware methodologies to improve the alignment between reasoning and the answer. Yet, those consistency problems have not yet been explored in the field of MER, where reasoning datasets are limited. This gap is significant, as MER applications require explanations that are not just plausible but emotionally coherent with the predicted affect. We therefore introduce a novel reward applied with GRPO training that explicitly enforces coherence between the generated explanation and the target emotion, addressing the unique challenge of achieving consistency in the absence of human-annotated rationales.
@@ 3 Method
Our primary objective is to develop an MLLM capable of emotion understanding that not only accurately predicts emotion labels from face video and speech audio but also generates consistent and faithful natural language explanations aligned with these predictions. Building upon recent advancements demonstrating the effectiveness of RL-based training for improving both emotion recognition accuracy and reasoning in MLLM (zhao2025r1), we integrate a novel Emotional Rationale Verifier (ERV) module into an established RL framework. This integration allows us to directly evaluate and enhance the emotional consistency of the generated explanations. Specifically, we extend the HumanOmni model architecture (zhao2025humanomni) and GRPO (guo2025deepseek), a reward-driven RL training strategy showing strong performance in various multimodal tasks (liu2025visualrft; chen2025r1v; zhao2025r1). While HumanOmni effectively unifies multimodal features and GRPO enhances CoT reasoning, existing approaches often struggle with emotionally incongruent explanations despite accurate label predictions, as highlighted in our introduction.
To precisely address this critical misalignment, our proposed ERV module assesses whether a generated explanation coherently supports the target emotion. Based on this assessment, we define a novel reward signal, the Explanation Reward ($R_{E}$), which encourages the MLLM to produce more faithful and emotion-aligned rationales. The overall training pipeline, detailing the ERV module and its integration into the GRPO framework, is visually represented in Figure 2.
@@ 3.1 Model Architecture and Training Overview
The training of this model follows a robust two-stage strategy: an initial Supervised Fine-Tuning (SFT) (zhang2024SFT) phase, followed by RL using GRPO. This progressive approach ensures the model first learns fundamental reasoning structures and then refines its explanation generation based on explicit reward signals.
Figure 2: (a): The policy model $\pi_{\theta}$ generates $G$ responses $o_{1}$, …, $o_{G}$, and the Emotional Rationale Verifier (ERV) assigns an explanation reward $R_{E}$ to each response. (b): For each response $o_{i}$, its explanation $E_{i}$ is extracted and evaluated to produce its reward $R_{i,E}$. Collectively, $R_{E}=\{R_{1,E},R_{2,E},\ldots,R_{G,E}\}$ denotes the set of explanation rewards for the $G$ responses.
Initializing Emotion Reasoning via SFT
The initial training phase involves SFT on a high-quality, albeit relatively small, dataset named EMER (lian2023EMER). This dataset consists of $(V,A)$ pairs annotated with emotion labels and detailed explanatory descriptions. The primary objective of this phase is to enable the MLLM to grasp the basic structural relationships between multimodal inputs, emotion labels, and their textual explanations. It specifically helps the model learn to produce outputs in the required format <<think>><</think>><<answer>><</answer>>. This SFT phase serves as a crucial initialization step for our policy model, denoted as $\pi_{\theta}$ (i.e., the trainable MLLM). During SFT, the model is optimized to maximize the likelihood of generating the target output sequence given the input. The training objective is the negative log-likelihood loss:
@@
Following the SFT phase, we transition to RL using GRPO to enhance the model’s reasoning capability and, critically, the consistency of its explanations. The SFT-trained model $\pi_{\theta}$ functions as the learnable policy, while a frozen duplicate, $\pi_{ref}$, acts as the reference model. For each input $(V,A,P)$, both the policy model and the reference model generate G candidate output sequences, $\{o_{1},o_{2},\dots,o_{G}\}$. Each generated sequence is then evaluated by a comprehensive reward function, which we detail in the next section. The policy $\pi_{\theta}$ is subsequently optimized to maximize the expected reward $R$, while simultaneously being regularized to remain close to the behavior of $\pi_{ref}$. The optimization objective of GRPO is as follows:
@@
While the GRPO framework is highly effective for reward-driven optimization, it traditionally lacks an explicit mechanism to guarantee the consistency between the predicted emotion and its generated explanation. To overcome this inherent limitation and directly address the core challenge identified in our introduction, we introduce the Emotional Rationale Verifier ($ERV$) and its associated reward ($R_{E}$) as a novel and crucial signal within the GRPO training loop.
@@ 3.2 Emotional Rationale Verifier (ERV): Design and Training
Although existing datasets such as EMER (lian2023EMER) and MERR-fine (cheng2024MERR) provide pairs of emotional descriptions and emotion labels, their limited sizes (332 and 4,487 samples, respectively) are insufficient for our setting. This is particularly problematic because the generated descriptions often contain not only emotional expressions but also contextual details about people, environments, and temporally evolving emotional states. In addition, the emotion label distribution is imbalanced. As a result, existing datasets lack the diversity and coverage necessary to effectively train the ERV.
To address the limitations of existing datasets in terms of both size and class imbalance, we construct a pseudo-labeled dataset consisting of emotional text descriptions paired with emotion labels. As shown in Figure 1, while R1-Omni (zhao2025r1) does not always produce perfectly emotion-aligned outputs, its generated descriptions are often emotionally plausible. Leveraging this property, we generate textual explanations from the training sets of DFEW (jiang2020dfew) and MAFW (liu2022mafw), and use a closed-source LLM (GPT-4.1) to assign up to two representative emotion labels to each description. This process yields a pseudo-labeled dataset of 20K emotional text–label pairs, which is used to train our ERV module. To further mitigate the class imbalance, we augment underrepresented categories, namely disgust, fear, contempt, disappointment, neutral, and helplessness, by generating additional emotional descriptions using GPT-4.1.
To train the ERV on the constructed dataset, we use the RoBERTa model trained on the GoEmotions dataset (demszky2020goemotions) [^1] https://huggingface.co/SamLowe/roberta-base-go˙emotions as a backbone. We fine-tune this model to predict emotion labels from text descriptions using our 23K-sample dataset, which includes both human-annotated and GPT-generated pseudo-labeled examples. The model predicts the emotion label based on the output of the [CLS] token, which encodes the global semantics of the input description. These predictions are then used to compute the Explanation Reward for each generated explanation $E_{i}$.